[The following information applies to the questions displayed below.) Jackson and Ashley Turner (both 45 years old) are marrib. How much can Ashley contribute if she files a separate return? Contribution to Roth IRAc. Assume that Ashley earned all of the couples income and that she contributed the maximum amount she is allowed to contrib
Answer #1

2020 Deduction Contribution Limit:

    - Lesser of $6,000 or earned income

    - For taxpayers 50 & above, lesser of $7,000 or earned income

Proportional phase-out of full limit of for MAGI (modified AGI) between $196,000 and $206,000


A) If they file jointly Ashley may contribute = $5,340


Proportional phase-out = (197,100 - 196,000) / (206,000 - 196,000) = 1,100 / 10,000 = 0.11 or 11%

11% of 6,000 = 660

Contribution = 6,000 - 660 = 5,340


B) If Ashley files a separate return she gets zero deductible contribution, because her MAGI is greater than $10,000 (phase-out limit).

Ashley's earned income is $98,550 = $197,100 / 2


C) Same as part A.


source: McGraw Hill's Taxation of Individuals Textbook Page 13 - 22,